OpenStudy (anonymous):

Help On This Question? Will Give Medal Nd Fan :) Pleaseee Help ^_^ Simplify: 4-c ----- 2c-8 ??

OpenStudy (anonymous):

@nincompoop @sarah_98 @Machida @malkassir @goformit100 @Eyad

OpenStudy (anonymous):

\[\frac{4-c}{2c-8}\] take from the denominator 2 as a common factor \[\frac{4-c}{2(c-4)}\] Is that familiar ? can you complete ?

OpenStudy (anonymous):

I Got the answer -1/2 from the book but iDont know wht to do after wht yu just did

OpenStudy (anonymous):

Hint : (4-c) If you take a -1 as a common factor it will be : -1(-4+c) which equals -1(c-4) .. Huh ? what do you think now ?

OpenStudy (anonymous):

@EsaaLoca ,lol ,you there ?

OpenStudy (anonymous):

Lol Yeeeh

OpenStudy (anonymous):

kewl,You got the point ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

\[\Huge \frac{-1\color{Red}{(c-4)}}{2\color{Red}{(c-4)}}\] Anything to cancel out ? anything annyyyything ? maybe the highlighted one ? :P

OpenStudy (anonymous):

\[\Huge Still~Vague~??\]

OpenStudy (anonymous):

oohhhhhhhhh!

OpenStudy (anonymous):

HolyCow !!!!

OpenStudy (anonymous):

Thank you C:

OpenStudy (anonymous):

Yvw :) Anytime @!